A box of mass 50 kg is at rest on a horizontal frictionless surface. A constant horizontal force F then acts on the box and accelerates it to the right. It is observed that it takes the box 6.9 seconds to travel 28 meters. What is the magnitude of the force?
Answer #1

Force = mass x acceleration, F = ma

m = 50 kg

Using the equation of motion, s = (u + v)/2 x t

we have u, s & t

s = 28m

t = 6.9 secs

u = 0

u + v = 28/(t/2)

therefore because u = 0, we find v

v = 28/(6.9/2)

v = 8.1159 m/s

we know that u = 0

a = (v - u)/t

= (8.1159 - 0)/6.9

a = 1.1762 m/s/s

F = 50 x 1.1762

F = 58.8 N